Anna and her 40 kg classmate, Maria is 3.00 x 10 -8 N.

Determine the mass of Anna if they are 2.0 m away from each other.


1
Expert's answer
2021-12-06T09:48:33-0500
"F=\dfrac{GM_AM_M}{d^2},""M_A=\dfrac{Fd^2}{GM_M},""M_A=\dfrac{3\times10^{-8}\ N\times(2\ m)^2}{6.67\times10^{-11}\ \dfrac{N\times m^2}{kg^2}\times40\ kg}=45\ kg."